Concurrency: Practice and Experience Status Report — July 7, 1999

(C332) “Distributed Context Switching: A Technique to Improve Turnaround Times of Distributed Computa-

tions” Authors: Donald McLaughlin and Parth Dasgupta Current Status: Referee comments to authors (9/17/97).

(C341) “Automatic Task-Level Parallelization with pTASK”

Authors: Tarek S. Abdelrahman, and Sum Huynk Current Status: Referee comments to authors (7/3/97).

(C348) “Dynamic Load Balancing for Object-Based Parallel computations”

Authors: Michele Di Santo, Franco Frattolillo, Wilma Russo, and Eugenio Zimeo Current Status: Referee comments to authors (9/17/97).

(C364) “PRAM Programming: In Theory and In Practice”

Authors: D. S. Lecomber, C. J. Siniolakis, and K. R. Sujithan Current Status: Referee comments to authors (10/27/97).

(C373) “Combining Linear Data Patterns for Accessing Parallel Memories through Artitrary Multistage Net-

works” Authors: Mayez Al-Mouhamed and Lubomir Bic Current Status: Referee comments to authors (1/22/98).

(C377) “Developing Parallel Applications in a Distributed Heterogeneous Environment”

Authors: Alvaro E. Campos, Miguel Nussbaum and Ricardo Schmutzer Current Status: Referee comments to authors (7/31/98).

(C379) “A Vector-based Strategy for Dynamic Resource Allocation”

Authors: Jerrell Watts, and Stephen Taylor Current Status: Referee comments to authors (7/29/98).

(C385) “The Design and Implementation of the Parallel Out-of-core ScaLAPACK LU, QR and Cholesky Fac-

torization Routines” Authors: Eduardo D’Azevedo, and Jack Dongarra Current Status: Referee comments to author (10/12/98).

(C387) “A Tool for Parallel Dynamic Programming Algorithms”

Authors: D. Gonzalez, F. Almeida, J. Roda, and C. Rodriguez Current Status: Referee comments to author (10/1/98).

(C392) “Synchronization Transformations for Parallel Computing”

Authors: Pedro C. Diniz and Martin C. Rinard Current Status: Referee comments to authors (3/25/99).

(C393) “After You, Alfonse and Gustave: A Toolkit for Developing Distributed Algorithms”

Authors: Steven Burdette and Tracy Camp Current Status: Referee comments to authors (3/25/99).

(C394) “Designing Shared-Memory Communication Devices for MPI Message-Passing Libraries” Authors: Boris V. Protopopov and Anthony Skjellum Current Status: Referee comments to authors (3/25/99).

(C395) “Object-Oriented Analysis and Design of the Message Passing System”

Authors: A. Skjellum, D. Wooley, Z. L, M. Wolf, P. Bangalore, A. Lumsdaine, J. Squyres, and B. McCandless Current Status: Referee comments to authors (3/25/99).

(C396) “On the Implementation of the Opus Coordination Language”

Authors: Erwin Laure, Matthew Haines, Piyush Mehrotra, and Hans Zima Current Status: Referee comments to authors (3/25/99).

(C398) “A Load Balanced Distributed Computing System”

Authors: M. Bozyigit, J. Al-ghamdi, H. Barada, and M. Ghouseuddin Current Status: Referee comments to authors (3/25/99).

(C402) “A Concurrent Process Creation Service to Support SPMD Based Parallel Processing on COWs” Authors: M. Hobbs, and A. Goscinski Current Status: Referee comments to authors (4/5/99).

(C403) “Theoretical Analysis of Optimal Striping Unit Size for Disk Array”

Authors: Hai Jin, and Kai Hwang Current Status: Letter to referees Alok Choudhary and Craig Patten (2/14/99).

(C404) “Integrating Task Parallel in Data Parallel Platforms for NOWs”

Authors: Binu K. J., and D-Janki Ram Current Status: Letter to referees G. Fagg and P. Mehrotra (4/8/99).

(C405) “A Parallel Language and Its Programming System for Heterogeneous Networks” Authors: A. Lastovetsky, D. Arapov, A. Kalinov, and I. Ledovskih Current Status: Letter to referees D. Walker and V. Sunderam (4/8/99).

(C407) “Concurrent Programing in VISO”

Authors: Mohammed Al-Mulhem Current Status: Letter to referees A. Hey and D. Lecomber (4/13/99).

(C408) “Extended Simaphore Operations”

Authors: S. Hodson, N. Dunstan, and I. Fris Current Status: Letter to referees P. B. Hansen, and G. Thiruvathukal (4/13/99).